# A Survey on Quantum Machine Learning: Basics, Current Trends, Challenges, Opportunities, and the Road Ahead

### v2

@@ II-F4 Error Mitigation
The NISQ era is characterized by relatively small-sized quantum circuits. Moreover, quantum noise affects state preparation, gate operations, and measurement. Due to the many qubits and large circuit depth required by QEC codes, it is impossible to implement QEC on NISQ devices. On the other hand, QEM techniques offer low-overhead solutions to implement quantum circuits in an accurate and reliable way [50][54].
II-F5 Fault-Tolerant Quantum Computing
@@
II-G Quantum Stack
Quantum computers operate under a fundamentally different set of rules than classical computers and can solve certain problems faster by exploiting quantum effects. The development stack in Figure 12 shows an abstraction of the quantum stages formulating a generalized quantum stack. Starting from the Hardware or qubit level, the abstractions go up to the application level composed of Quantum Algorithms. Since QC is at an early development stage, industries primarily focus on the Quantum Hardware level, where the reliability of measurements plays a crucial role in designing and choosing the physical mediums. While simulated quantum hardware can deal with perfect qubits, realistic quantum hardware can be implemented with superconducting qubits, ion trapped qubits, quantum dots, and neutral atom qubits [59].
@@ II-G Quantum Stack
III Benefits of Quantum Computing
@@
III-A Quantum Dimensionality Reduction
A common way to reveal properties of an unknown quantum state, having many copies of a system in that state, is to measure different observables and to analyze the results statistically. The unknown quantum state can play an active role in this analysis. Given multiple copies of a quantum system with its respective density matrix, it is possible to perform the unitary transformation. Its result generates quantum coherence among different copies of the system. Popular techniques for dimensionality reduction are the Principal Component Analysis (PCA) [60], where the eigenvectors corresponding to the unknown state’s eigenvalues state are revealed in a very fast manner, and the Quantum Linear Discriminant Analysis Dimensionality Reduction (QLDADR) [61], where the vectors having a high-dimensional feature space are projected onto a low-dimensional feature space.
III-B Quantum Supremacy for Solving Complex Problems
The inherent problems’ difficulty, or hardness, is a fundamental concept in computational complexity theory (see Appendix A for more details). The hardness quantifies the resources required by different models of computation to solve problems, such as the number of steps of a deterministic Turing machine [62]. Various models and resources are often considered, including deterministic, nondeterministic, and probabilistic models, time and space constraints, and interactions across different models. “Quantum Supremacy” is a concept that indicates that a programmable quantum device can solve a problem that no classical computer can solve in a feasible amount of time. For instance, Google’s Sycamore processor [63] performs in $200$ seconds a task that would require $10,000$ years using a classical computer. Such an acclaimed advantage of QC compared to classical computers demands a modification of the complexity theory for problems that are solvable on a quantum computer. Hence, the classical complexity classification diagram has been redefined to incorporate the quantum perspective and evaluate the complexity of the issues, considering their ability to be solvable on a quantum computer. Figure 13 shows that a set of problems outside the P space that can be solved in Bounded-error, Quantum, and Polynomial-time (BQP) [64]. Identifying BQP problems is crucial to prove the overwhelming ability of quantum computers compared to classical computers.
@@ III-B Quantum Supremacy for Solving Complex Problems
III-C Quantum Enhancement to Classical Problems
@@
Demonstrating such aspects of QC is essential to gain confidence and value in using such systems. Hence, works like [65] that aim to practically demonstrate enhancements like quantum speedup for a computational problem are good examples of paths leading to practical quantum usability. Similar works investigate how the generic quantum advantage passing criteria can be tuned slightly towards strict reasoning to generate useful results and efforts.
To be critical of the current state of quantum practicality, existing methods are either heuristic-based or depend largely on the encoding of data in a quantum mechanical way that causes the change in performance and are the sole aspects making the difference from classical performance. In light of the hybrid quantum-classical systems, the work in [66] discusses important aspects that should be changed to obtain meaningful developments and revolves around how heuristic-based algorithms do not provide formal evidence that showcases their genuine and consistent advantage over classical algorithms. Moreover, the variational circuits can only implement linear classifiers on input quantum mechanically encoded to improve feature extraction, which can be replaced by classical support vector machines if the encoding is classically tractable.
@@
III-D Preserving Maximum Information
III-E Large-Scale Infrastructures and Design Automation
@@
The realization of large-scale quantum systems will take several years. However, it is the ultimate target for all significant quantum hardware companies. Their roadmaps aim to have such large-scale systems available for industrial and practical use around the end of the NISQ era. Figure 14 shows the development roadmaps provided by IBM [67] and Honeywell [68] for their current progress and plans towards achieving large-scale quantum infrastructures. These milestones demonstrate the continuous focus on developing quantum chips with more qubits every year, like the 432-qubit Osprey machine by IBM [69]. Similarly, Honeywell’s H-series [70] has stages planned from H1 to H5 as their progress from 10 qubit systems to larger scale qubit quantum systems envisioned to be designed around the year 2030.
IV Quantum Machine Learning Algorithms
IV-A Categorization of QML Approaches
Before diving into the details of QML algorithms, it is important to characterize different approaches based on the type of data and type of processor used to solve the problem [8]. The four categories (see Figure 16) are formed based on whether the data is classical (C) or quantum (Q) and whether the algorithm runs on a classical (C) or quantum (Q) computer.
@@ IV-A Categorization of QML Approaches
IV-B Parametrized Quantum Circuits
Variational or Parametrized Quantum Circuits (PQCs) are specific types of quantum algorithms that depend on free parameters. PQCs allow us to utilize the existing quantum computers to their full extent. In the context of QML, PQCs are used either to encode the data, where the parameters are determined by the data being encoded, or as a quantum model, where the parameters are determined by an optimization process. PQCs can be interpreted as ML models, considering a variational quantum classifier that uses two variational circuits (see Figure 17). The first circuit associates the gate parameters with fixed data inputs, while the second circuit depends on free and trainable parameters. This setup provides the basic building blocks to build QML algorithms on NISQ devices [71].
Figure 17: Example of a parametrized quantum circuit.
A Variational Quantum Algorithm (VQA) combines a classical optimizer with a PQC. VQAs represent a promising approach to achieving quantum advantage over classical systems, even when using NISQ devices that come with various constraints. These constraints include a limited number of qubits, restricted qubit connectivity, and decoherence errors that limit the depth of quantum circuits [72]. Despite these challenges, VQAs can leverage the strengths of both classical and quantum computing to tackle problems that are intractable for classical algorithms alone. Examples of applications in which VCAs excel are quantum chemistry, material science, financial modeling, and cryptography.
IV-C Quantum Optimization Problems
An optimization problem aims to find the best solution to a given challenge. For instance, the goals of a business are to minimize the production cost or maximize its revenue. The solution can either be discrete if the variable to optimize is determined from a countable set or continuous if the optimized value is found from a continuous function. It can also be either single-objective or multi-objective. Applying QML to solve optimization problems enables us to employ quantum classification and regression programs on QML infrastructure (see Figure 18). The intrinsic parallelism of quantum computing can speed up the optimization computation to compute the global minimum (or maximum) faster than classical computing [73].
@@ IV-C Quantum Optimization Problems
The primary goal and advantage of all efforts behind quantum algorithms is to motivate and identify the practical advantage over classical in practical scenarios. Similarly, in gate-based quantum computers, which are currently adopted in the industry, and in shared future directions, it is important and useful to define beneficial use-cases of a quantum algorithm, to showcase its efficient utilization of the quantum hardware. Such efforts are now making their way into the research work involving techniques like Filtering Variational Quantum Eigensolvers [74]. Driven by the idea of filtering operators while processing to increase speed and reliability towards converging towards an optimal solution, such efforts are necessary to make QML systems impactful.
IV-C1 Variational Amplitude Amplification and Estimation
@@
Amplitude amplification is a technique in quantum computing algorithms derived by generalizing the working principle of Grover’s Search Algorithm [75]. Initially discovered by Brassad and Hoyer [76], the approach to finding a solution using “amplitude amplification” gives rise to a new family of quantum algorithms with the advantage of obtaining a quadratic speedup over several classical algorithms in comparison to a brute-force search. However, to observe its best potential impact when applied to real-world problems that embody larger search spaces for solutions, we need heuristics along with ”amplitude amplification” to make it practically useful [77].
@@
Quantum Amplitude Estimation (QAE) is the task of finding an estimate for the amplitude of the solution state using the process of amplitude amplification described above. It can be generalized as a way of finding a solution to a combinatorial search problem, where components of interest in the space are projected in sub-spaces as a superposition of the required solution and other possibilities. Given the problem in this context, the process of “amplitude estimation” is the algorithm for estimating the probability that the final measurement after applying amplification will yield the desired state of the solution for the respective problem. The advantage of using this algorithm lies in the fact that it provides a quadratically faster solution than expected by a classical algorithm, provided the condition that the solution to the problem is known to be unique [75, 78, 77]. The first implementation of the algorithm by Brassad and Hoyer [76] for amplitude estimation employs a strategy involving phase estimation, which makes the algorithm compute-intensive. Another variant employs an iterative strategy to avoid phase estimation, which represents a useful evolution of the amplitude estimation strategy [79].
IV-C2 Quantum Approximate Optimization Algorithm
@@
The Quantum Approximate Optimization Algorithm (QAOA) [80] is a technique that finds approximate solutions to combinatorial optimization problems. It is based on PQCs that approximate the adiabatic evolution from an initial Hamiltonian. A set of parameters are tweaked to optimize a cost function out of the quantum circuit output. The QAOA method is a promising candidate for achieving quantum advantage over classical systems in the NISQ era. Potential applications of the QAOA in the real world span from the logistics field, such as designing air/ground traffic and shipping routes, to the finance domain, where it can be used to maximize profits and minimize risks for a given portfolio.
IV-C3 Variational Quantum Eigensolver
@@
A Variational Quantum Eigensolver (VQE) is a hybrid algorithm that uses both classical computers and quantum computers to find the ground state of a given physical system. The VQE algorithm is versatile and can be applied to a wide range of tasks, making it a promising candidate for use in many applications of quantum computers, including machine learning and control theory [81]. One particularly useful application of the VQE algorithm is finding the ground state energy (i.e., the minimum eigenvalue) of a Hamiltonian. This is accomplished by minimizing a cost function based on the expectation value of the Hamiltonian. An additional term in this algorithm accounts for the overlap between the excited and ground states.
@@
An instance of VQE requires the definition of two algorithmic sub-components, which are a quantum trial state (also called Ansatz [82]) and a classical optimizer. The optimizer varies the parameters of the Anzatz, such that it works towards a state, as determined by its parameters, that results in the minimum expectation value being measured by the input operator (Hamiltonian). Additionally, the VQE can be used to model complex wave functions in polynomial time, making it one of the most promising NISQ applications for quantum computing. In practical applications, the VQE is used in chemistry for simulations of molecules, as well as logistics and network design.
IV-C4 Quantum Annealing
The Quantum Annealing (QA) method is an algorithm to solve combinatorial optimization problems. Instead of using the temperature to explore the problem space, the QA uses the laws of quantum mechanics to measure the energy state. Starting from the qubits’ dual state, where all the possible combinations of the values are equally likely to happen, the quantum mechanical effects are gradually reduced through a process called quantum fluctuation [83] to reach the most optimal point where the lowest energy state is achieved. In practical use-cases, the QA is efficient for combinatorial optimization problems whose search space has several local minima, such as the travelling salesman problem.
@@ IV-C4 Quantum Annealing
QA techniques require an objective function to be defined in the Ising model or Quadratic Unconstrained Binary Optimisation (QUBO). The Ising model is a mathematical model in statistical mechanics [84]. The QUBO [85] is mathematically equivalent to the Ising model and can be used to formulate a problem in a more simple way.
In physics, the minimum energy principle [86] states that the internal energy decreases and approaches the minimum values. If we can formulate our problems as an energy minimization problem, the QA method can search for the best possible solution by utilizing quantum methods over an energy landscape.
IV-D Quantum Neural Networks
Quantum Neural Networks (QNNs) are computational Artificial Neural Network (ANN) models that are based on the principles of quantum mechanics [87]. As shown in Figure 20, typically in QNNs, the input data is loaded with classical data inputs. The quantum circuit contains a feature map module with input parameters and an Ansatz module with trainable weights. Measurements are conducted to obtain the outputs.

### v3

@@ 2.6.4 Error Mitigation
Despite sharing the same goal as QEC to reduce the impact of quantum noise, Quantum Error Mitigation (QEM) techniques operate differently. While QEC aims to restore the correct value after the error occurs, QEM aims to reduce or suppress the errors that occur during computation. Moreover, as shown in Figure 11, QEC completely removes the noise, while QEM keeps a small amount of noise under control. A scenario where QEC is applied, called ‘hard regime’, is theoretically ideal but difficult to realize. The other scenario where QEM is employed, called ‘easy regime’, is experimentally easier to implement but at the price of lower qubit protection. For QEM deployment, Mitiq (LaRose et al., 2022) is an open-source error mitigation package in Python that implements QEM techniques like zero-noise extrapolation (Giurgica-Tiron et al., 2020) and probabilistic error cancellation (van den Berg et al., 2023) for quantum computers.
The NISQ era is characterized by relatively small-sized quantum circuits. Moreover, quantum noise affects state preparation, gate operations, and measurement. Due to the many qubits and large circuit depth required by QEC codes, it is impossible to implement QEC on NISQ devices. On the other hand, QEM techniques offer low-overhead solutions to implement quantum circuits in an accurate and reliable way (Shaib et al., 2021; Nielsen & Chuang, 2010a).
2.6.5 Fault-Tolerant Quantum Computing
@@
2.7 Quantum Stack
Quantum computers operate under a fundamentally different set of rules than classical computers and can solve certain problems faster by exploiting quantum effects. The development stack in Figure 12 shows an abstraction of the quantum stages formulating a generalized quantum stack. Starting from the Hardware or qubit level, the abstractions go up to the application level composed of Quantum Algorithms. Since QC is at an early development stage, industries primarily focus on the Quantum Hardware level, where the reliability of measurements plays a crucial role in designing and choosing the physical mediums. While simulated quantum hardware can deal with perfect qubits, realistic quantum hardware can be implemented with superconducting qubits, ion trapped qubits, quantum dots, and neutral atom qubits (Bertels et al., 2021).
3 Benefits of Quantum Computing
@@
3.1 Quantum Dimensionality Reduction
A common way to reveal properties of an unknown quantum state, having many copies of a system in that state, is to measure different observables and to analyze the results statistically. The unknown quantum state can play an active role in this analysis. Given multiple copies of a quantum system with its respective density matrix, it is possible to perform the unitary transformation. Its result generates quantum coherence among different copies of the system. Popular techniques for dimensionality reduction are the Principal Component Analysis (PCA) (Lloyd et al., 2014), where the eigenvectors corresponding to the unknown state’s eigenvalues state are revealed in a very fast manner, and the Quantum Linear Discriminant Analysis Dimensionality Reduction (QLDADR) (Yu et al., 2023), where the vectors having a high-dimensional feature space are projected onto a low-dimensional feature space.
3.2 Quantum Supremacy for Solving Complex Problems
@@
The inherent problems’ difficulty, or hardness, is a fundamental concept in computational complexity theory (see Appendix A for more details). The hardness quantifies the resources required by different models of computation to solve problems, such as the number of steps of a deterministic Turing machine (Papadimitriou & Yannakakis, 1991). Various models and resources are often considered, including deterministic, nondeterministic, and probabilistic models, time and space constraints, and interactions across different models. “Quantum Supremacy” is a concept that indicates that a programmable quantum device can solve a problem that no classical computer can solve in a feasible amount of time. For instance, Google’s Sycamore processor (Frank Arute, 2019) performs in $200$ seconds a task that would require $10,000$ years using a classical computer. Such an acclaimed advantage of QC compared to classical computers demands a modification of the complexity theory for problems that are solvable on a quantum computer. Hence, the classical complexity classification diagram has been redefined to incorporate the quantum perspective and evaluate the complexity of the issues, considering their ability to be solvable on a quantum computer. Figure 13 shows that a set of problems outside the P space that can be solved in Bounded-error, Quantum, and Polynomial-time (BQP) (Watrous, 2009). Identifying BQP problems is crucial to prove the overwhelming ability of quantum computers compared to classical computers.
3.3 Quantum Enhancement to Classical Problems
@@
Demonstrating such aspects of QC is essential to gain confidence and value in using such systems. Hence, works like Pokharel & Lidar (2023) that aim to practically demonstrate enhancements like quantum speedup for a computational problem are good examples of paths leading to practical quantum usability. Similar works investigate how the generic quantum advantage passing criteria can be tuned slightly towards strict reasoning to generate useful results and efforts.
To be critical of the current state of quantum practicality, existing methods are either heuristic-based or depend largely on the encoding of data in a quantum mechanical way that causes the change in performance and are the sole aspects making the difference from classical performance. In light of the hybrid quantum-classical systems, the work of Liu et al. (2021) discusses important aspects that should be changed to obtain meaningful developments and revolves around how heuristic-based algorithms do not provide formal evidence that showcases their genuine and consistent advantage over classical algorithms. The work of Kashif et al. (2024) demonstrates that HQNNs exhibit better scalability than classical neural networks when increasing classification task complexity. Moreover, the variational circuits can only implement linear classifiers on input quantum mechanically encoded to improve feature extraction, which can be replaced by classical support vector machines if the encoding is classically tractable.
3.4 Preserving Maximum Information
3.5 Large-Scale Infrastructures and Design Automation
@@
The realization of large-scale quantum systems will take several years. However, it is the ultimate target for all significant quantum hardware companies. Their roadmaps aim to have such large-scale systems available for industrial and practical use around the end of the NISQ era. Figure 14 shows the development roadmaps provided by IBM (Quantum, 2022) and Honeywell (Honeywell, 2022a) for their current progress and plans towards achieving large-scale quantum infrastructures. These milestones demonstrate the continuous focus on developing quantum chips with more qubits every year, like the 432-qubit Osprey machine by IBM (IBM, 2022). Similarly, Honeywell’s H-series (Honeywell, 2022b) has stages planned from H1 to H5 as their progress from 10 qubit systems to larger scale qubit quantum systems envisioned to be designed around the year 2030.
4 Quantum Machine Learning Algorithms
@@
4.1 Categorization of QML Approaches
Before diving into the details of QML algorithms, it is important to characterize different approaches based on the type of data and type of processor used to solve the problem (Aïmeur et al., 2006). The four categories (see Figure 17) are formed based on whether the data is classical (C) or quantum (Q) and whether the algorithm runs on a classical (C) or quantum (Q) computer.
@@ 4.1 Categorization of QML Approaches
4.2 Parametrized Quantum Circuits
Variational or Parametrized Quantum Circuits (PQCs) are specific types of quantum algorithms that depend on free parameters. PQCs allow us to utilize the existing quantum computers to their full extent. In the context of QML, PQCs are used either to encode the data, where the parameters are determined by the data being encoded, or as a quantum model, where the parameters are determined by an optimization process. PQCs can be interpreted as ML models, considering a variational quantum classifier that uses two variational circuits (see Figure 17). The first circuit associates the gate parameters with fixed data inputs, while the second circuit depends on free and trainable parameters. This setup provides the basic building blocks to build QML algorithms on NISQ devices (Benedetti et al., 2019).
A Variational Quantum Algorithm (VQA) combines a classical optimizer with a PQC. VQAs represent a promising approach to achieving quantum advantage over classical systems, even when using NISQ devices that come with various constraints. These constraints include a limited number of qubits, restricted qubit connectivity, and decoherence errors that limit the depth of quantum circuits (Cerezo et al., 2021). Despite these challenges, VQAs can leverage the strengths of both classical and quantum computing to tackle problems that are intractable for classical algorithms alone. Examples of applications in which VCAs excel are quantum chemistry, material science, financial modeling, and cryptography.
4.3 Quantum Optimization Problems
An optimization problem aims to find the best solution to a given challenge. For instance, the goals of a business are to minimize the production cost or maximize its revenue. The solution can either be discrete if the variable to optimize is determined from a countable set or continuous if the optimized value is found from a continuous function. It can also be either single-objective or multi-objective. Applying QML to solve optimization problems enables us to employ quantum classification and regression programs on QML infrastructure (see Figure 18). The intrinsic parallelism of quantum computing can speed up the optimization computation to compute the global minimum (or maximum) faster than classical computing (Dasari et al., 2019).
@@ 4.3 Quantum Optimization Problems
The primary goal and advantage of all efforts behind quantum algorithms is to motivate and identify the practical advantage over classical in practical scenarios. Similarly, in gate-based quantum computers, which are currently adopted in the industry, and in shared future directions, it is important and useful to define beneficial use-cases of a quantum algorithm, to showcase its efficient utilization of the quantum hardware. Such efforts are now making their way into the research work involving techniques like Filtering Variational Quantum Eigensolvers (Amaro et al., 2022). Driven by the idea of filtering operators while processing to increase speed and reliability towards converging towards an optimal solution, such efforts are necessary to make QML systems impactful.
4.3.1 Variational Amplitude Amplification and Estimation
@@
Amplitude amplification is a technique in quantum computing algorithms derived by generalizing the working principle of Grover’s Search Algorithm (Grover, 1996). Initially discovered in Brassard & Hoyer , the approach to finding a solution using “amplitude amplification” gives rise to a new family of quantum algorithms with the advantage of obtaining a quadratic speedup over several classical algorithms in comparison to a brute-force search. However, to observe its best potential impact when applied to real-world problems that embody larger search spaces for solutions, we need heuristics along with "amplitude amplification" to make it practically useful (Brassard et al., 2002).
@@
Quantum Amplitude Estimation (QAE) is the task of finding an estimate for the amplitude of the solution state using the process of amplitude amplification described above. It can be generalized as a way of finding a solution to a combinatorial search problem, where components of interest in the space are projected in sub-spaces as a superposition of the required solution and other possibilities. Given the problem in this context, the process of “amplitude estimation” is the algorithm for estimating the probability that the final measurement after applying amplification will yield the desired state of the solution for the respective problem. The advantage of using this algorithm lies in the fact that it provides a quadratically faster solution than expected by a classical algorithm, provided the condition that the solution to the problem is known to be unique (Grover, 1996; 1997; Brassard et al., 2002). The first implementation of the algorithm in Brassard & Hoyer for amplitude estimation employs a strategy involving phase estimation, which makes the algorithm compute-intensive. Another variant employs an iterative strategy to avoid phase estimation, which represents a useful evolution of the amplitude estimation strategy (Grinko et al., 2021).
4.3.2 Quantum Approximate Optimization Algorithm
@@
The Quantum Approximate Optimization Algorithm (QAOA) (Farhi et al., 2014) is a technique that finds approximate solutions to combinatorial optimization problems. It is based on PQCs that approximate the adiabatic evolution from an initial Hamiltonian. A set of parameters are tweaked to optimize a cost function out of the quantum circuit output. The QAOA method is a promising candidate for achieving quantum advantage over classical systems in the NISQ era. Potential applications of the QAOA in the real world span from the logistics field, such as designing air/ground traffic and shipping routes, to the finance domain, where it can be used to maximize profits and minimize risks for a given portfolio.
4.3.3 Variational Quantum Eigensolver
@@
A Variational Quantum Eigensolver (VQE) is a hybrid algorithm that uses both classical computers and quantum computers to find the ground state of a given physical system. The VQE algorithm is versatile and can be applied to a wide range of tasks, making it a promising candidate for use in many applications of quantum computers, including machine learning and control theory (Tilly et al., 2022). One particularly useful application of the VQE algorithm is finding the ground state energy (i.e., the minimum eigenvalue) of a Hamiltonian. This is accomplished by minimizing a cost function based on the Hamiltonian’s expectation value. An additional term in this algorithm accounts for the overlap between the excited and ground states.
@@
An instance of VQE requires the definition of two algorithmic sub-components, which are a quantum trial state (also called Ansatz (Zhang, 2012)) and a classical optimizer. The optimizer varies the parameters of the Anzatz, such that it works towards a state, as determined by its parameters, that results in the minimum expectation value being measured by the input operator (Hamiltonian). Additionally, the VQE can be used to model complex wave functions in polynomial time, making it one of the most promising NISQ applications for quantum computing. In practical applications, the VQE is used in chemistry for simulations of molecules, as well as logistics and network design.
4.3.4 Quantum Annealing
@@
The Quantum Annealing (QA) method is an algorithm to solve combinatorial optimization problems. Instead of using the temperature to explore the problem space, the QA uses the laws of quantum mechanics to measure the energy state. Starting from the qubits’ dual state, where all the possible combinations of the values are equally likely to happen, the quantum mechanical effects are gradually reduced through a process called quantum fluctuation (Park & Nha, 2023) to reach the most optimal point where the lowest energy state is achieved (see Figure 19). In practical use-cases, the QA is efficient for combinatorial optimization problems whose search space has several local minima, such as the traveling salesman problem.
QA techniques require an objective function to be defined in the Ising model or Quadratic Unconstrained Binary Optimisation (QUBO). The Ising model is a mathematical model in statistical mechanics (Borle & Lomonaco, 2019). The QUBO (Jun, 2022) is mathematically equivalent to the Ising model and can be used to formulate a problem in a more simple way.
In physics, the minimum energy principle (Callen, 1985) states that the internal energy decreases and approaches the minimum values. If we can formulate our problems as an energy minimization problem, the QA method can search for the best possible solution by utilizing quantum methods over an energy landscape.
4.4 Quantum Neural Networks